Use standard heats of formation (delta H) and standard entropies (delta S)to calculate Go for the following reaction at 25oC. Which type of reaction does it fall under? Is it spontaneous? Does it have a temperature (delta G) would equal zero?

NH3(g) ---->H2(g)+ N2(g)
Answer #1

NH3(g) ---->H2(g)+ N2(g)

delta H = delta Hfproducts - delta Hfreactants

           = (0 + 0) - (- 45.9)

         = 45.9 kJ /mol

delta S = (191.6 + 130.7) - (192.8)

           = 129.5 J / K

delta G = delta H - T delta S

           = 45.9 - 298 x 0.1295

delta G = 7.309 kJ/mol

this is endothermic reaction .

this is not spontaneous. because delta G > 0

T = delta H / delta S

    = 45.9 x 10^3 / 129.5

T = 354 K